How much acid and water are needed to dilute 12.0 M concentrated hydrochloric acid HCl, to make 4.00 liters of a 3.00 M solution?


1
Expert's answer
2022-06-08T04:17:03-0400

1.

Using the formula n = CM * V we find how many moles of HCl are in 4 liters of 3 M solution:

n = 3 mol/l * 4 l = 12 mol HCl

2.

We find the volume of a 12 M solution containing 12 mol of HCl:

V = n / CM = 12mol / 12mol/l = 1 liters

Answer: 1 liters 12M HCl solution and

3 liters H2O (water)